Can you help me solve this -2(2t+8)+5=5 ; -4(5t-2)+2t-8=54 ; 5(3c+5)-4c=3c-7 and last 6i+3(i-8)=2(2i+8)? I really just want to KNOW THE STEPS TO FIND THE VARIABLE NOT just the answer. Thank you!
Answer by lynnlo(4176)

(1)t=-4
(2)t=-3
(3)c=-4
(4)i=8 example:-2(2t+8)+5=5
-2(8+2t)+5=5
(8*-2+2t*-2)+5=5
rewrite (-16+5=-11
combine like terms -16+5=-11
-11+-4t=5
solve -11t-4t=5
move all terms containing t to the left side,everything else to the right side
add-11+11t-4t=5+11
combine
-11+11=0
0t-4t=5+11
-4t=16
divide each side by -4
t=-4